# TabXEval: Why this is a Bad Table? An eXhaustive Rubric for Table Evaluation

### v1

@@ Abstract
Evaluating tables qualitatively & quantitatively presents a significant challenge, as traditional metrics often fail to capture nuanced structural and content discrepancies. To address this, we introduce a novel, methodical rubric integrating multi-level structural descriptors with fine-grained contextual quantification, thereby establishing a robust foundation for comprehensive table comparison. Building on this foundation, we propose TabXEval, an eXhaustive and eXplainable two-phase evaluation framework. TabXEval initially aligns reference tables structurally via TabAlign & subsequently conducts a systematic semantic and syntactic comparison using TabCompare; this approach clarifies the evaluation process and pinpoints subtle discrepancies overlooked by conventional methods. The efficacy of this framework is assessed using TabXBench, a novel, diverse, multi-domain benchmark we developed, featuring realistic table perturbations and human-annotated assessments. Finally, a systematic analysis of existing evaluation methods through sensitivity-specificity trade-offs demonstrates the qualitative and quantitative effectiveness of TabXEval across diverse table-related tasks and domains, paving the way for future innovations in explainable table evaluation.
TabXEval: Why this is a Bad Table?
An eXhaustive Rubric for Table Evaluation
thanks: These authors contributed equally to this work.Vihang Pancholi [^1] footnotemark: 1Jainit Bafna211footnotemark: 1Tejas Anvekar1 Manish Shrivastava2thanks: Primary supervisor of this work.Vivek Gupta1 1Arizona State University  2IIIT Hyderabad {vpancho1,tanvekar,vgupta140}@asu.edu {jainit.bafna,m.shrivastava}@research.iiit.ac.in
@@ 1 Introduction
Tables are a ubiquitous data format across a multitude of domains from financial reporting and medical data analysis to scientific research and knowledge bases. Each of these fields confronts its own unique challenges: financial tables must maintain rigorous numerical precision and consistency, while medical tables demand exactitude in categorical and temporal data; both scenarios underscore the difficulty of comparing tables when discrepancies may arise from subtle semantic, structural, or formatting variations. Traditional evaluation metrics originally designed for text generation, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), and chrF Popović (2015), rely on surface-level n–gram overlaps and fail to capture these multifaceted discrepancies. Even embedding-based approaches like BERTScore Zhang et al. (2020), which excel in measuring semantic similarity in text-to-text tasks, fall short when applied to the intricacies of tabular data.
Recent efforts have sought to address these limitations. Benchmarks such as StructBenchGu et al. (2024) and TanQAkhtar et al. (2025) introduce complex structural challenges and query-oriented tasks that push the boundaries of table generation. In parallel, methods like the “Is this a bad table?” metric Ramu and Bandyopadhyay (2024) decompose tables into atomic natural language statements to assess errors in a more granular fashion. However, these approaches often compromise either on the explainability or the granularity needed for robust table evaluation. Moreover, in contrast to domains like image captioning where datasets and benchmarks such as THumB Kasai et al. (2022) provide comprehensive human evaluations to validate metrics table evaluation lacks a comparable level of benchmark diversity and human-aligned assessment.
@@
To overcome these challenges, we propose a novel evaluation framework, TabXEval, which is built upon a comprehensive, multi-level rubric that integrates structural descriptors with fine-grained contextual quantification. TabXEval operates in two distinct phases: TabAlign first establishes a precise baseline alignment of table elements via both deterministic and LLM-based methods, and TabCompare subsequently conducts a detailed semantic and syntactic analysis of aligned cells. This two-phase approach enables us to capture both coarse (table-level) and fine-grained (cell-level) discrepancies that conventional metrics overlook.
Complementing our evaluation method is TabXBench, a synthetic yet diverse benchmark designed to mimic real-world table perturbations across multiple domains. By incorporating human-annotated assessments aligned with our proposed rubric, TabXBench serves as a crucial testbed for quantifying sensitivity, specificity, and overall alignment with human judgment addressing a significant gap in current evaluation paradigms.
@@
  • We demonstrate TabXEval’s qualitative and quantitative effectiveness in text-to-table generation, fostering explainable automatic evaluation.
Our TabXBench dataset and TabXEval source code are available at https://coral-lab-asu.github.io/tabxeval/.
2 Related Works
Early work in text-to-table generation predominantly leveraged datasets originally designed for table-to-text tasks such as RotoWireWiseman et al. (2017), E2ENovikova et al. (2017), WikiBioLebret et al. (2016), and WikiTableTextPasupat and Liang (2015) which, while pioneering, suffer from inherent limitations. These datasets are often characterized by a lack of structural diversity and a tendency to produce hallucinated responses, thereby compromising both the factual accuracy and robustness of generated tables. Recent benchmarks, including StrucBenchGu et al. (2024)and TanQAkhtar et al. (2025), have emerged to address these shortcomings by introducing more complex structural challenges and query-oriented generation tasks that emphasize multi-hop reasoning and multi-source data integration. However, these new benchmarks remain under-explored, underscoring the need for further work to validate their effectiveness and bridge the gap between theoretical advancements and practical performance.
Traditional evaluation metrics for generated text such as BLEU Papineni et al. (2002), ROUGELin (2004), METEORBanerjee and Lavie (2005), and chrFPopović (2015) primarily assess surface-level n–gram overlaps between candidate and reference texts. To capture deeper semantic similarity, embedding-based methods like BERTScoreZhang et al. (2020) were developed; yet, when applied to table evaluation, they fall short. In the context of tables, key structural elements such as row and column alignment, header consistency, and hierarchical relationships are critical. Metrics that operate on a token or cell–level basis, including Exact Match, PARENTDhingra et al. (2019), and even BERTScoreZhang et al. (2020), fail to account for the interdependent and contextual nature of table cells, leading to an incomplete assessment of both semantic and structural fidelity.
Recent work in StructBench Gu et al. (2024)introduces the H-score and p-score for table evaluation. The H-score measures hierarchical alignment and contextual consistency, while the p-score incorporates probabilistic confidence intervals to assess prediction reliability. These metrics address the limitations of token-based methods by focusing on structural coherence and robustness.
To overcome these limitations, recent research has explored hybrid approaches that combine structural decomposition with semantic evaluation. For example, the “Is this a bad table?” metric Ramu and Bandyopadhyay (2024) decomposes tables into natural language atomic statements and applies entailment-based scoring to capture both syntactic and semantic discrepancies in a holistic manner. Together, these efforts highlight an ongoing shift towards more exhaustive and explainable evaluation frameworks such as our proposed TabXEval that jointly consider both the structural alignment and contextual integrity of tables, thereby paving the way for more robust and generalizable table-to-table comparisons.
@@ 3 TabXEval
We establish a transparent evaluation protocol for reference table evaluation. Our rubrics and rules are developed through rigorous research survey and designed to increase the reliability of evaluation.
@@ 3.1 Evaluation Rubric
To ensure consistent and fair structural evaluation, we introduce a comprehensive rubric that promotes clarity and transparency in reference-based table assessment. It quantifies correctness and information coverage through four key evaluation categories:
Structure Descriptor provides a coarse assessment at the table level capturing missing, extra, and matched information to reflect overall integrity.
Column Descriptor detects column data types and highlights missing or extra content. This enables fine-grained evaluation, especially for heterogeneous columns (e.g., dates, numbers, strings), avoiding uniform penalization.
Cell-Level Descriptor evaluates both semantic and syntactic correctness of cell values. Unlike prior methods Zhang et al. (2020); Ramu and Bandyopadhyay (2024); Gu et al. (2024), it correctly handles variant representations of the same data, reducing false penalties.
Granular Cell-Level Difference measures the degree of mismatch at the cell level, accounting for format shifts (e.g., meters to cm, years to full dates), and captures absolute differences crucial for fine-grained scoring.
3.2 Method
@@
This formulation captures the multi-level nature of table evaluation insipred by proposed rubric. First aggregating errors across different information types (Missing, Extra, Partial) via the outer summation, and then evaluating the correctness at various entity levels (row, column, cell) using the inner summation. The term $\gamma_{p}$ further refines the score by quantifying discrepancies in partially matched cells through a normalized absolute difference between the ground truth ($GT$) and the reference ($Ref$), ensuring that both coarse structural errors and fine-grained content differences are robustly accounted for in a single, interpretable metric.
@@ 3.3 TabXBench
Dataset # of Tables # Perturbation/Table Average
@@ 4 Experiments
LLMs. Throughout our experiments, we used GPT-4o, Gemini 2.0-flash, and LLaMA-3.3-Instruct 70B. All models were executed with identical sampling settings (default temperature, top-$k$, and top-$p$) unless specified otherwise. All our prompts for TabAlign, TabCompare and Direct-LLM basline are given in Figure 6, Figure 7, Figure 8.
@@ 4.1 Human correlation
For each sample in the ground truth set of TabXBench, two human evaluators evaluated the tables using our proposed rubrics and guidelines. For both Human and TabXEval, we present ground-truth and a randomly selected perturbation (out of 5) to fill the rubrics. Finally these rubrics are compared using both Pearson Sedgwick (2012) and Kendall’s Tau Sen (1968) correlation coefficients to quantify the degree of alignment between our method and human ratings. We observe very high correlation, i.e, $\mathbf{99.7\%}$ and $\mathbf{95.1\%}$ Pearson’s $\rho$ correlation for the Rubric Structure Descriptor and Cell Level Descriptor respectively. In contrast, Direct LLM based baseline correlates, only $\mathbf{30.6\%}$ and $\mathbf{40.6\%}$ Pearson’s $\rho$ correlation for the Rubric Structure Descriptor and Cell Level Descriptor respectively. Revealing that it fails to understand the rubric, and quantify the structural and contextual challenges in table evaluation, and hence is unable to correctly align with humans correlation. Similarly we report $\mathbf{99.1\%}$ and $\mathbf{92.8\%}$ human correlation using Kendall’s $\tau$. While the baseline is $\mathbf{30.7\%}$ and $\mathbf{55\%}$ on Structure Descriptor and Cell Level Descriptor rubric respectively. These results demonstrate that disentangling the alignment and comparison phases is critical for robust table evaluation as Direct method falls short even hen presented with evaluation rubrics. TabXEval’s correlation with human are consistent in capturing multi-level nuances for real-world challenges in table evaluation resembling humans.
@@ 4.2 Human Ranking Correlation Study
Metrics Spearman’s Kendall’s W-Kendall’s RBO Spearman’s
$\rho$uparrow $\tau$uparrow $\tau^{\dagger}$uparrow $\uparrow$ Footrule $\downarrow$
@@
Figure 3: Sensitivity–Specificity trade-off across metrics. Bubble size reflects the harmonic mean of accuracy, sensitivity, and specificity; color shows F1-score. The dashed line marks ideal balance—top-right methods perform best. Green denotes optimal (Goldilocks) zone; pink favors sensitivity, yellow favors specificity.
@@ 4.3 What Sets TabXEval Apart?
Comparisons with Other Metrics.
  • P-Score sits near the Goldilocks zone, indicating solid table-level correctness. However, it lacks explainability and granular insights, offering only a single 0010101010 score limiting interpretability and error traceability.
  • H-Score and BERTScore better capture semantics than string-based metrics (EM, ROUGE-L, chrF), but often miss structural errors like swapped columns or missing rows leading to moderate sensitivity but poor specificity.
  • TabEval relies on entailment over LLM-generated atomic statements, often missing fine-grained numeric/unit mismatches. It showed false positives on re-formatted but semantically correct tables in our tests.
@@
Table 3: Cell-Level Performance Analysis of Extra (EI), Missing (MI), and Partial mismatches across data types — numerical, string, boolean, date, list, time, and other — for WikiTables, WikiBio, TANQ, and RotoWire. Highlights: GPT-4o shows fewer string EI in WikiTables and lower partial errors in numerical and string cells in TANQ and RotoWire.
4.4 Performance Analysis
Adaptability and Robustness Evaluation Figure 4 illustrates that TabXEval consistently outperforms existing metrics in human-correlation across a broad range of weighting schemes. Each box plot corresponds to 26 permutations of weights for key dimensions (e.g., missing, extra, row, column, cell, partial) these weights contribute to the Equations 3.2 &  3.2, set to either perm(0)=$\{0,1\}$ or perm=(0.25)$\{0.25,1\}$. This flexibility lets TabXEval adapt to dataset-specific priorities (e.g., penalizing missing rows more heavily) or maintain robust, domain-agnostic performance. Notably, our best-performing configuration (marked by the star) demonstrates both high average correlation and low variance, underscoring TabXEval’s capacity to balance fine-grained descriptors with overall structural fidelity.
5 Conclusion
6 Limitations
7 Ethics Statement
Acknowledgements
We thank the Complex Data Analysis and Reasoning Lab at Arizona State University for computational support, and Krishna Singh Rajput for his support during the submission process.
Appendix
Appendix A Performance Analysis
@@
Method Spearman’s $\rho$ Kendall’s $\tau$ Weighted Kendall’s $\tau$ RBO Spearman’s Footrule
@@
TabXEval (LLaMA3.3 70B-Instruct) 0.37 0.30 0.30 0.29 0.44
TabXEval (Qwen72B-Instruct) 0.33 0.27 0.28 0.30 0.38
TabXEval (Gemini 2.0 pro) 0.30 0.23 0.21 0.30 0.40
@@
Table 5: Robustness of TabXEval across LLM back-bones. Human-ranking correlations for TabXEval run with GPT-4o, LLaMA-3.3-Instruct, Qwen-72B-Instruct and Gemini-2.0-pro, compared to BLEURT and the P-Score baseline. TabXEval + GPT-4o attains the strongest alignment (Spearman’s $\rho$ = 0.44), but all back-bones retain a $\geq$ 0.30 correlation, indicating backbone-agnostic reliability.
Figure 5: Human Ranking Correlation, existing metric vs ours all configurations.
Appendix B Examples

### v2

@@ Abstract
Evaluating tables qualitatively and quantitatively poses a significant challenge, as standard metrics often overlook subtle structural and content-level discrepancies. To address this, we propose a rubric-based evaluation framework that integrates multi-level structural descriptors with fine-grained contextual signals, enabling more precise and consistent table comparison. Building on this, we introduce TabXEval, an eXhaustive and eXplainable two-phase evaluation framework. TabXEval first aligns reference and predicted tables structurally via TabAlign, then performs semantic and syntactic comparison using TabCompare, offering interpretable and granular feedback. We evaluate TabXEval on TabXBench, a diverse, multi-domain benchmark featuring realistic table perturbations and human annotations. A sensitivity-specificity analysis further demonstrates the robustness and explainability of TabXEval across varied table tasks. Code and data are available at https://coral-lab-asu.github.io/tabxeval/.
TabXEval: Why this is a Bad Table? An eXhaustive Rubric for
Table Evaluation
thanks: These authors contributed equally to this work.Vihang Pancholi[Uncaptioned image]11footnotemark: 1Jainit Bafna[Uncaptioned image]11footnotemark: 1Tejas Anvekar[Uncaptioned image] Manish Shrivastava[Uncaptioned image]thanks: Primary supervisor and corresponding author of this work.Vivek Gupta[Uncaptioned image] [Uncaptioned image] Arizona State University  [Uncaptioned image] IIIT Hyderabad {vpancho1,tanvekar,vgupta140}@asu.edu {jainit.bafna,m.shrivastava}@research.iiit.ac.in
Figure 1: Sensitivity–Specificity trade-off across metrics. Bubble size reflects the harmonic mean of accuracy, sensitivity, and specificity; color shows F1-score. The dashed line marks ideal balance top-right methods perform best. Green denotes optimal (Goldilocks) zone; pink favors sensitivity, yellow favors specificity.
@@ 1 Introduction
Tables are a ubiquitous data format across critical workflows: budget forecasts, patient dashboards, and experimental logs alike where even a one-cell error can trigger costly re-statements or clinical misinterpretations. As large language models (LLMs) and other neural systems are increasingly tasked with generating or transforming such tables, reliable automatic evaluation becomes a bottleneck.
Despite the structured nature of tables, most evaluation metrics treat them as plain text. Metrics like BLEU, ROUGE, METEOR, and chrF rely on n-gram overlap, ignoring row–column alignment and unit consistency Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005); Popović (2015). Embedding-based scores such as BERTScore improve semantic sensitivity but overlook structural errors like column swaps Zhang* et al. (2020). Token-level metrics, including Exact Match and PARENT, address factual grounding but fail under reordered or merged schemas Dhingra et al. (2019). Structural benchmarks highlight these issues: StructBench exposes failures on partial cell mismatches Gu et al. (2024), TanQ reveals brittleness under unit conversions Akhtar et al. (2025), and Data-QuestEval sacrifices structure for corpus-level QA-based comparisons Rebuffel et al. (2021). Atomic decomposition methods like “Is this a bad table?” improve detection but add opacity and computational cost Ramu et al. (2024). In contrast, work like THumB shows the benefit of rubric-based human ratings in improving evaluation transparency Kasai et al. (2022).
Taken together, existing metrics tend to emphasize either semantics or structure, but rarely both. They offer limited diagnostic insight, often masking specific error types and failing to provide actionable feedback for model improvement. As shown in image-captioning work like THumB, coupling automatic scores with rubric-based human ratings yields more interpretable and reliable evaluations Kasai et al. (2022). These findings underscore the need for a rubric-based evaluation framework that explicitly assesses both structural alignment and semantic fidelity. In contrast to single-score metrics that collapse diverse errors—such as schema mismatches, contextual omissions, or subtle content shifts—into a single value, a rubric-based approach offers fine-grained, interpretable feedback. Such granularity is essential for complex or high-stakes tasks, where even minor discrepancies can significantly affect downstream performance.
@@
To overcome these challenges, we introduce TabXEval, a novel evaluation framework built on a structured, multi-level rubric that combines high-level structural descriptors with fine-grained contextual signals. TabXEval operates in two phases: TabAlign first performs precise alignment of table elements using both rule-based and LLM-assisted strategies, followed by TabCompare, which conducts detailed semantic and syntactic analysis over the aligned cells. This design allows TabXEval to capture both table-level and cell-level discrepancies that prior metrics often overlook.
To rigorously test our rubric and framework, we construct TabXBench, a diverse, synthetic benchmark that emulates realistic table perturbations across multiple domains. TabXBench includes human-annotated ratings grounded in our rubric, serving as a gold standard for evaluating metric sensitivity, specificity, and alignment with human judgment. By providing controlled, interpretable scenarios, TabXBench fills a critical gap in current evaluation practice, enabling robust and explainable assessment of structured table outputs. Unlike prior metrics, TabXEval, as shown in Figure 1, excels at detecting subtle discrepancies i.e. sensitive and accurately localizing errors between tables i.e. specific enough. We summarize our main contributions below:
@@
  • We present TabXEval’s qualitative and quantitative effectiveness in table generation task, enabling explainable automatic evaluation.
@@ 2 TabXEval
We establish a transparent and systematic protocol for table evaluation by comparing a reference table—the candidate table produced by a human or an LLM with a ground-truth table. These tables may differ in formatting, interpretation, or unit representation, necessitating a rigorous evaluation framework. For instance, an LLM-generated table might omit entire rows or abbreviate numeric values (e.g., “100k” vs. “100,000”), while a human-curated table may specify units only in the header (e.g., “velocity” vs. “velocity (m/s)”). To address such discrepancies, we design a set of rules based rubrics TabXEval, which aim to improve the reliability and interpretability of table evaluation.
@@ 2.1 Evaluation Rubric
Towards ensuring consistency and fairness in structure evaluation we provide an exhaustive rubric that enhances clarity and transparency in reference-based table evaluation. To quantify the degree of correctness and coverage of information at the most granular level, we advocate 4 categories of evaluation protocols:
Structure Descriptor coarsely assesses the overall structure against ground truth. This component compares information at the table level (missing information, extra information and exact matches), giving a high level description of information integrity.
Column Descriptor identifies data types of each column based on missing and extra information. This allows for a fine-grained strategy for evaluating cell values, as tables with heterogeneous columns like dates, numbers, strings etc should not be penalized on the same scale.
Cell Level Descriptor looks at both semantic and syntactical representation of cell values. Other methods such as Zhang* et al. (2020); Ramu et al. (2024); Gu et al. (2024) fail in recognizing different representations of the same data and hence can wrongly penalize correct information. This component of the rubric is necessary to craft an ideal table evaluation framework.
Granular Cell Level Difference that determines the magnitude of discrepancies between reference and ground truth table necessary to quantify instances w.r.t. cell level descriptions. It also captures variations by changing the format(e.g. m to cm, years to date) to report the absolute differences.
2.2 TabXEval Rubric
@@
This formulation captures the multi-level nature of table evaluation insipred by proposed rubric. First aggregating errors across different information types (Missing, Extra, Partial) via the outer summation, and then evaluating the correctness at various entity levels (row, column, cell) using the inner summation. The term $\gamma_{p}$ further refines the score by quantifying discrepancies in partially matched cells through a normalized absolute difference between the ground truth ($GT$) and the reference ($Ref$), ensuring that both coarse structural errors and fine-grained content differences are robustly accounted for in a single, interpretable metric. An illustrative example demonstrating the application of the above equations is provided in Appendix B.
@@ 2.3 TabXBench Benchmark
Dataset # of Tables # Perturb/Table Average
@@ 3 Experiments
LLMs. Throughout our experiments, we used GPT-4o, Gemini 2.0-flash, and LLaMA-3.3-Instruct 70B. All models were executed with identical sampling settings (default temperature, top-$k$, and top-$p$) unless specified otherwise. All our prompts for TabAlign, TabCompare and Direct-LLM basline are given in Appendix C.
@@ 3.1 Human correlation
For each sample in the ground truth set of TabXBench, two human evaluators evaluated the tables using our proposed rubrics and guidelines. For both Human and TabXEval, we present ground-truth and a randomly selected perturbation (out of 5) to fill the rubrics. The detailed human annotation protocol is described in Appendix F. Finally these rubrics are compared using both Pearson Sedgwick (2012) and Kendall’s Tau Sen (1968) correlation coefficients to quantify the degree of alignment between our method and human ratings. We observe very high correlation, i.e, $\mathbf{99.7\%}$ and $\mathbf{95.1\%}$ Pearson’s $\rho$ correlation for the Rubric Structure Descriptor and Cell Level Descriptor respectively.
In contrast, Direct LLM based baseline correlates, only $\mathbf{30.6\%}$ and $\mathbf{40.6\%}$ Pearson’s $\rho$ correlation for the Rubric Structure Descriptor and Cell Level Descriptor respectively. Revealing that it fails to understand the rubric, and quantify the structural and contextual challenges in table evaluation, and hence is unable to correctly align with humans correlation. Similarly we report $\mathbf{99.1\%}$ and $\mathbf{92.8\%}$ human correlation using Kendall’s $\tau$. While the baseline is $\mathbf{30.7\%}$ and $\mathbf{55\%}$ on Structure Descriptor and Cell Level Descriptor rubric respectively. These results demonstrate that disentangling the alignment and comparison phases is critical for robust table evaluation as Direct method falls short even hen presented with evaluation rubrics. TabXEval’s correlation with human are consistent in capturing multi-level nuances for real-world challenges in table evaluation resembling humans.
@@ 3.2 Human Ranking Correlation Study
Metrics Spearman’s $\rho$uparrow Kendall’s $\tau$uparrow W-Kendall’s $\tau^{\dagger}$uparrow RBO $\uparrow$ Spearman’s Footrule $\downarrow$
@@ 3.3 What Sets TabXEval Apart?
Comparisons with Other Metrics. We compare TabXEval against several widely used metrics: (1) P-Score performs well at a high level and sits near the Goldilocks zone in our evaluations, reflecting strong table-level correctness. However, it lacks explainability and granular insights, offering only a single 0010101010 score that limits interpretability and error traceability. (2) H-Score and BERTScore better capture semantics than string-based metrics such as EM, ROUGE-L, and chrF, but often overlook structural errors like swapped columns or missing rows—resulting in moderate sensitivity but poor specificity. (3) TabEval uses entailment over LLM-generated atomic statements, but frequently misses fine-grained numeric or unit mismatches. In our experiments, it produced false positives on tables that were re-formatted yet semantically equivalent.
@@
Table 3: Cell-Level Performance Analysis of Extra (EI), Missing (MI), and Partial mismatches across data types numerical, string, boolean, date, list, time, and other for WikiTables, WikiBio, TANQ, and RotoWire. Highlights: GPT-4o shows fewer string EI in WikiTables and lower partial errors in numerical and string cells in TANQ and RotoWire.
@@
3.4 Performance Analysis
Metrics Spearman’s $\rho$uparrow Kendall’s $\tau$uparrow W-Kendall’s $\tau^{\dagger}$uparrow RBO $\uparrow$ Spearman’s Footrule $\downarrow$
@@ 3.4 Performance Analysis
TabXEval (LLaMA) 0.37 0.30 0.30 0.29 0.44
TabXEval (Qwen) 0.33 0.27 0.28 0.30 0.38
TabXEval (Gemini) 0.30 0.23 0.21 0.30 0.40
@@
Table 5: Robustness of TabXEval across LLM back-bones. Human-ranking correlations for TabXEval run with GPT-4o, LLaMA-3.3 70B-Instruct, Qwen-72B-Instruct and Gemini-2.0-pro, compared to BLEURT and the P-Score baseline. TabXEval + GPT-4o attains the strongest alignment (Spearman’s $\rho$ = 0.44), but all back-bones retain a $\geq$ 0.30 correlation, indicating backbone-agnostic reliability.
Adaptability and Robustness Evaluation Figure 4 illustrates that TabXEval consistently outperforms existing metrics in human-correlation across a broad range of weighting schemes (Appendix A Figure 5). Each box plot corresponds to 26 permutations of weights for key dimensions (e.g., missing, extra, row, column, cell, partial); these weights contribute to Equations 2.2 and 2.2, set either as perm(0)=$\{0,1\}$ or perm(0.25)=$\{0.25,1\}$. This flexibility allows TabXEval to adapt to dataset-specific priorities (e.g., penalizing missing rows more heavily) while maintaining robust, domain-agnostic performance. Notably, our best-performing configuration (marked by the star) demonstrates both high average correlation and low variance, underscoring TabXEval’s capacity to balance fine-grained descriptors with overall structural fidelity.
Complementing this, Table 5 demonstrates TabXEval’s backbone-agnostic robustness: it achieves strong alignment with human rankings across diverse LLMs, including GPT-4o, LLaMA, Qwen, and Gemini. Even under varying model architectures, TabXEval maintains a correlation of $\geq$ 0.30 across all metrics, reinforcing its reliability beyond just weighting flexibility. A detailed qualitative analysis, provided in Appendix D, further demonstrates TabXEval’s effectiveness across both domain-agnostic and specific weighting configurations.
4 Comparison with Related Work
Text-to-Table Generation.
Early text–to–table research exploited single–domain corpora such as RotoWire for basketball summaries Wiseman et al. (2017), the E2E restaurant set Novikova et al. (2017), WikiBio infobox–biography pairs Lebret et al. (2016), and WikiTableText Pasupat and Liang (2015). While pioneering, these resources offer limited schema variety and often encourage hallucinated or under-structured outputs. Recent collections address these gaps: StructBench permutes headers, merges columns, and shuffles schemas to test structural generalisation Gu et al. (2024), whereas TanQ requires multi-hop, multi-source reasoning to generate answer tables Akhtar et al. (2025). Such challenging benchmarks expose systematic weaknesses in both generation models and legacy evaluation metrics, motivating the fine-grained rubric employed by TabX.
Other Evaluation Metrics
1. Surface and embedding overlap. Classic n-gram scores BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), and chrF Popović (2015) treat a table as flat text, ignoring header alignment or cell hierarchy. Embedding-based BERTScore Zhang* et al. (2020) improves semantic sensitivity, yet still overlooks structural fidelity. Token-level Exact Match and PARENT Dhingra et al. (2019) partially reward factual grounding, but cannot detect column swaps or unit shifts.
2. Structure-aware and reference-less scores. StructBench introduces H-Score and P-Score, targeting hierarchical integrity and holistic quality, respectively Gu et al. (2024). TabEval (“Is this a bad table?”) decomposes each table into atomic statements and uses textual entailment to capture fine-grained errors Ramu et al. (2024). Complementarily, Data-QuestEval dispenses with references altogether by generating and answering questions directly over the source data, achieving strong human correlation in data-to-text tasks Rebuffel et al. (2021). We further show that our approach performs robustly across a wide range of table structures and can be extended to handle hierarchical formats, as discussed in Section E.
Despite these advances, existing metrics still emphases either semantics or structure and provide a single numeric values with limited error traceability or explainability. Our two-phase TabX closes this gap by disentangling alignment (TabAlign) from cell-level comparison (TabCompare), producing an interpretable score that balances sensitivity and specificity.
5 Conclusion and Future Work
We found two key directions for future work. Firstly, while our current approach leverages large language models to ensure robustness and generalization to unseen table structures, their computational overhead can hinder practical deployment. Developing more compact alternatives using smaller models (e.g., BART, T5) would require large-scale, heterogeneous fine-tuning data to preserve performance on out-of-distribution tables—a resource that is currently limited. Future efforts could explore model distillation or semi-supervised training to create lightweight yet reliable variants of TabXEval. Secondly, although TabXEval effectively handles structural variations—such as merged or decomposed cells—through LLM-guided alignment, directly supporting complex hierarchical tables (e.g., nested cells or multi-level headers) remains a challenge. A promising direction is to incorporate a preprocessing step that flattens hierarchical structures into standardized key paths (e.g., transforming “Release Date” with sub-headers “Month” and “Year” into “Release Date.Month” and “Release Date.Year”), enabling compatibility with our approach while preserving semantic structure.
Limitations
Ethics Statement
Acknowledgements
We thank the Complex Data Analysis and Reasoning Lab at Arizona State University for computational support, and Krishna Singh Rajput for his support during the submission process. Lastly, we thank our lab cat, Coco, for her unwavering commitment to keeping our professor creatively unhinged during deadline season.
Appendix
Appendix A Performance Analysis
Figure 5 represent human ranking correlation of TabXEval across various confirgrations of parameters.
Figure 5: Human Ranking Correlation, ours all configurations.
Appendix B Illustration Example